OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = –67?

OpenStudy (callisto):

a1 = a17 - 16d 13 = -67 - 16d d =? 35th term = a1 + (n-1)d = 13 + (35-1) d = 13 + (34) (?) = ...

OpenStudy (anonymous):

-157

OpenStudy (callisto):

Yes